# Supersonic OT: Fast Unconditionally Secure Oblivious Transfer

### v1

@@ 2.1 Notations
By $\epsilon$ we mean an empty string. We denote a sender by ${S}$ and a receiver by ${R}$. We assume parties interact with each other through a regular secure channel. $U$ denotes a universe of messages $m_{\scriptscriptstyle 1},\ldots,m_{\scriptscriptstyle t}$. We define $\sigma$ as the maximum size of messages in $U$, i.e., $\sigma=Max(|m_{\scriptscriptstyle 1}|,\ldots,|m_{\scriptscriptstyle t}|)$.
@@ 2.2 Security Model
$\{\mathsf{Sim}_{\scriptscriptstyle 1}(x,f_{\scriptscriptstyle 1}(x,y))\}_{\scriptscriptstyle x,y}\stackrel{{\scriptstyle c}}{{\equiv}}\{\mathsf{View}_{\scriptscriptstyle 1}^{\scriptscriptstyle\Gamma}(x,y)\}_{\scriptscriptstyle x,y}$
$\{\mathsf{Sim}_{\scriptscriptstyle 2}(y,f_{\scriptscriptstyle 2}(x,y))\}_{\scriptscriptstyle x,y}\stackrel{{\scriptstyle c}}{{\equiv}}\{\mathsf{View}_{\scriptscriptstyle 2}^{\scriptscriptstyle\Gamma}(x,y)\}_{\scriptscriptstyle x,y}$
@@ 4.1 Security Definition
Let $\mathcal{F}_{\scriptscriptstyle\mathcal{OT}^{\scriptscriptstyle 2}_{\scriptscriptstyle 1}}$ be the OT functionality defined above. We assert that protocol $\Gamma$ realizes $\mathcal{F}_{\scriptscriptstyle\mathcal{OT}^{\scriptscriptstyle 2}_{\scriptscriptstyle 1}}$ in the presence of a passive adversary, if for every non-uniform PPT adversary $\mathcal{A}$ in the real model, there is a non-uniform PPT simulator $\mathtt{Sim}$ in the ideal model, where:
$\begin{split}\Big{\{}\mathtt{Sim}_{\scriptscriptstyle{S}}\big{(}(m_{\scriptscriptstyle 0},m_{\scriptscriptstyle 1}),\epsilon\big{)}\Big{\}}_{\scriptscriptstyle m_{0},m_{1},s}\stackrel{{\scriptstyle c}}{{\equiv}}\Big{\{}\mathtt{View}_{\scriptscriptstyle{S}}^{\scriptscriptstyle\Gamma}\big{(}(m_{\scriptscriptstyle 0},m_{\scriptscriptstyle 1}),\epsilon,s\big{)}\Big{\}}_{\scriptscriptstyle m_{0},m_{1},s}\end{split}$ (1)
$\begin{split}\Big{\{}\mathtt{Sim}_{\scriptscriptstyle{P}}(\epsilon,\epsilon)\Big{\}}_{\scriptscriptstyle m_{0},m_{1},s}\stackrel{{\scriptstyle c}}{{\equiv}}\Big{\{}\mathtt{View}_{\scriptscriptstyle{P}}^{\scriptscriptstyle\Gamma}\big{(}(m_{\scriptscriptstyle 0},m_{\scriptscriptstyle 1}),\epsilon,s\big{)}\Big{\}}_{\scriptscriptstyle m_{0},m_{1},s}\end{split}$ (2)
$\begin{split}\Big{\{}\mathtt{Sim}_{\scriptscriptstyle{R}}\Big{(}s,\mathcal{F}_{\scriptscriptstyle\mathcal{OT}^{\scriptscriptstyle 2}_{\scriptscriptstyle 1}}\big{(}(m_{\scriptscriptstyle 0},m_{\scriptscriptstyle 1}),\epsilon,s\big{)}\Big{)}\Big{\}}_{\scriptscriptstyle m_{0},m_{1},s}\stackrel{{\scriptstyle c}}{{\equiv}}\Big{\{}\mathtt{View}_{\scriptscriptstyle{R}}^{\scriptscriptstyle\Gamma}\big{(}(m_{\scriptscriptstyle 0},m_{\scriptscriptstyle 1}),\epsilon,s\big{)}\Big{\}}_{\scriptscriptstyle m_{0},m_{1},s}\end{split}$ (3)
@@ 4.3 Proof of Security
Also, in the real model, the pair $e^{\prime}$ that is given to ${P}$ is always permuted based on the value of ${S}$’s share (i.e., $s_{\scriptscriptstyle 1}\in\{0,1\}$) which is not known to ${P}$; whereas, in the ideal model, the pair $v$ is not permuted. However, given the permuted pair $e^{\prime}$ and not permuted pair $v$, a distinguisher cannot tell where each pair has been permuted with a probability greater than $\frac{1}{2}$. Therefore, Relation 2 (in Section 4.1) holds.

### v2

@@ 2.1 Notations
By $\epsilon$ we mean an empty string. We denote a sender by ${S}$ and a receiver by ${R}$. We assume parties interact with each other through a regular secure channel. By the notation, $\mathcal{X}\stackrel{{\scriptstyle c}}{{\equiv}}\mathcal{Y}$, we mean that the two distributions $\mathcal{X}$ and $\mathcal{Y}$ are computationally indistinguishable. By $\mathcal{X}{\equiv}\mathcal{Y}$ we mean $\mathcal{X}$ and $\mathcal{Y}$ are unconditionally indistinguishable. $U$ denotes a universe of messages $m_{\scriptscriptstyle 1},\ldots,m_{\scriptscriptstyle t}$. We define $\sigma$ as the maximum size of messages in $U$, i.e., $\sigma=Max(|m_{\scriptscriptstyle 1}|,\ldots,|m_{\scriptscriptstyle t}|)$.
@@ 2.2 Security Model
$\{\mathsf{Sim}_{\scriptscriptstyle 1}(x,f_{\scriptscriptstyle 1}(x,y))\}_{\scriptscriptstyle x,y}{\equiv}\{\mathsf{View}_{\scriptscriptstyle 1}^{\scriptscriptstyle\Gamma}(x,y)\}_{\scriptscriptstyle x,y}$
$\{\mathsf{Sim}_{\scriptscriptstyle 2}(y,f_{\scriptscriptstyle 2}(x,y))\}_{\scriptscriptstyle x,y}{\equiv}\{\mathsf{View}_{\scriptscriptstyle 2}^{\scriptscriptstyle\Gamma}(x,y)\}_{\scriptscriptstyle x,y}$
[…]
@@ 4.1 Security Definition
Let $\mathcal{F}_{\scriptscriptstyle\mathcal{OT}^{\scriptscriptstyle 2}_{\scriptscriptstyle 1}}$ be the OT functionality defined above. We assert that protocol $\Gamma$ realizes $\mathcal{F}_{\scriptscriptstyle\mathcal{OT}^{\scriptscriptstyle 2}_{\scriptscriptstyle 1}}$ in the presence of a passive adversary, if for every adversary $\mathcal{A}$ in the real model, there is a simulator $\mathtt{Sim}$ in the ideal model, where:
$\begin{split}\Big{\{}\mathtt{Sim}_{\scriptscriptstyle{S}}\big{(}(m_{\scriptscriptstyle 0},m_{\scriptscriptstyle 1}),\epsilon\big{)}\Big{\}}_{\scriptscriptstyle m_{0},m_{1},s}{\equiv}\Big{\{}\mathtt{View}_{\scriptscriptstyle{S}}^{\scriptscriptstyle\Gamma}\big{(}(m_{\scriptscriptstyle 0},m_{\scriptscriptstyle 1}),\epsilon,s\big{)}\Big{\}}_{\scriptscriptstyle m_{0},m_{1},s}\end{split}$ (1)
$\begin{split}\Big{\{}\mathtt{Sim}_{\scriptscriptstyle{P}}(\epsilon,\epsilon)\Big{\}}_{\scriptscriptstyle m_{0},m_{1},s}{\equiv}\Big{\{}\mathtt{View}_{\scriptscriptstyle{P}}^{\scriptscriptstyle\Gamma}\big{(}(m_{\scriptscriptstyle 0},m_{\scriptscriptstyle 1}),\epsilon,s\big{)}\Big{\}}_{\scriptscriptstyle m_{0},m_{1},s}\end{split}$ (2)
$\begin{split}\Big{\{}\mathtt{Sim}_{\scriptscriptstyle{R}}\Big{(}s,\mathcal{F}_{\scriptscriptstyle\mathcal{OT}^{\scriptscriptstyle 2}_{\scriptscriptstyle 1}}\big{(}(m_{\scriptscriptstyle 0},m_{\scriptscriptstyle 1}),\epsilon,s\big{)}\Big{)}\Big{\}}_{\scriptscriptstyle m_{0},m_{1},s}{\equiv}\Big{\{}\mathtt{View}_{\scriptscriptstyle{R}}^{\scriptscriptstyle\Gamma}\big{(}(m_{\scriptscriptstyle 0},m_{\scriptscriptstyle 1}),\epsilon,s\big{)}\Big{\}}_{\scriptscriptstyle m_{0},m_{1},s}\end{split}$ (3)
@@ 4.3 Proof of Security
Also, in the real model, the pair $e^{\prime}$ that is given to ${P}$ is always permuted based on the value of ${S}$’s share (i.e., $s_{\scriptscriptstyle 1}\in\{0,1\}$) which is not known to ${P}$; whereas, in the ideal model, the pair $v$ is not permuted. However, given the permuted pair $e^{\prime}$ and not permuted pair $v$, a distinguisher cannot tell where each pair has been permuted with a probability greater than $\frac{1}{2}$. Therefore, Relation 2 (in Section 4.1) holds. $\Box$